# Black hole interiors of homogeneous holographic solids under shear strain

### v2

@@ 4.2 Emergence of domain wall geometry under large shear deformations
where $f_{0}=\frac{M\nu^{3}}{(3+M\nu)(3+\nu^{2})}$, $u_{*}$ satisfies $m^{2}\left(\frac{e^{\Omega}}{2}\right)^{M}u_{*}^{6M/\nu}=\frac{9}{3+M\nu}$ and $\chi_{h}\equiv\chi(u_{h})=\int^{u_{h}}_{0}\frac{1}{2}uh^{\prime}(u)^{2}du$. Under sufficiently large shear deformation, the analytical solution (26) accurately describes the bulk geometry from the vicinity of the UV boundary, through the event horizon, and deep into the interior. Near the UV boundary $u=u_{*}$, it exhibits an exact scaling symmetry—the anisotropic fixed point identified in Ref. Baggioli:2020qdg . More interestingly, the solution (26) also remains valid in the region behind the event horizon, extending to a scale $u_{\text{K}}\gg u_{h}$ that is pushed toward the singularity as the shear deformation increases. At $u=u_{\text{K}}$, the geometry exhibits a Kasner scaling symmetry, whose features will be detailed in the next section. As will be shown later, $u=u_{\text{K}}$ marks the transition point between two Kasner geometries. In the large shear limit $\Omega\rightarrow\infty$, with $u_{\text{K}}$ approaching the singularity, this analytic solution dominates the entire bulk spacetime.

### v3

@@ 4.2 Emergence of domain wall geometry under large shear deformations
with $\nu=\frac{3M}{M+2N}$ and $f_{0}=\frac{M\nu^{3}}{(3+M\nu)(3+\nu^{2})}$. $u_{*}$ satisfies $m^{2}\left(\frac{e^{\Omega}}{2}\right)^{M}u_{*}^{6M/\nu}=\frac{9}{3+M\nu}$ and $\chi_{h}\equiv\chi(u_{h})=\int^{u_{h}}_{0}\frac{1}{2}uh^{\prime}(u)^{2}du$. Under sufficiently large shear deformation, the analytical solution (24) accurately describes the bulk geometry from the vicinity of the UV boundary, through the event horizon, and deep into the interior. Near the UV boundary $u=u_{*}$, it exhibits an exact scaling symmetry—the anisotropic fixed point identified in Ref. Baggioli:2020qdg . More interestingly, the solution (24) also remains valid in the region behind the event horizon, extending to a scale $u_{\text{K}}\gg u_{h}$ that is pushed toward the singularity as the shear deformation increases. At $u=u_{\text{K}}$, the geometry exhibits a Kasner scaling symmetry, whose features will be detailed in the next section. As will be shown later, $u=u_{\text{K}}$ marks the transition point between two Kasner geometries. In the large shear limit $\Omega\rightarrow\infty$, with $u_{\text{K}}$ approaching the singularity, this analytic solution dominates the entire bulk spacetime.
@@ 1 $N\geq 1$ $\frac{M+2N}{M}$ $\frac{-M^{2}+(M+2N)^{2}}{3M^{2}+(M+2N)^{2}}$
To close this section, we draw a comparison between the Kasner transition in our model and that predicted by the standard BKL dynamics with pure gravitational walls. In the latter scenario, the exponents of the subsequent Kasner epoch are uniquely fixed by those of the preceding epoch via the relation (3.14) derived in Belinsky:1970ew . In our model, the exponents of the Kasner epoch induced by large shear deformation depend exclusively on potential form. If the transition were to adhere to the usual BKL dynamics of Belinsky:1970ew , the subsequent Kasner epoch would be fixed and unique. However, as is clearly illustrated in the right panels of Fig. 6 and Fig. 9), the subsequent Kasner epoch is not fixed. Instead, it depends sensitively on both the magnitude of the shear strain and the temperature. We therefore conclude that our results go beyond the transition rule of the standard BKL dynamics.
@@ Appendix A On the constraint $V_{X}\geq 0$
Next, we examine the physical interpretation of the scalar background profile $\bar{\phi}^{I}(x^{i})\sim x^{i}$ and how the form of the potential $V$ governs the mechanism that breaks translations at the boundary. Approaching the boundary, the gravity sector and matter fields become increasingly decoupled from each other. In this regime, the linearized equation for $\delta\phi^{I}$ is given by
$\partial_{\mu}\left[\sqrt{-\bar{g}}\left(\bar{V}_{X}\partial^{\mu}\delta\phi^{I}+\bar{V}_{Z}\bar{Z}\bar{\mathcal{I}}_{IJ}\partial^{\mu}\delta\phi^{J}\right)\right]=0,$ (42)
where $\mathcal{I}_{IJ}$ is the inverse of $\mathcal{I}^{IJ}$ defined in Eq. (2). Suppose that $V(\bar{X},\bar{Z})\sim u^{A}$ with the exponent $A>0$ near the boundary. Substituting this scaling into Eq. (42), the scalar fluctuations exhibit the near-boundary behavior
$\delta\phi^{I}(u,t,x^{I})\approx\delta\phi^{I}_{(0)}(t,x^{I})+\dots+\delta\phi^{I}_{(1)}(t,x^{I})\,u^{5-A}+\dots,$ (43)
where we have used $\bar{X}\sim u^{2}$ and $\bar{Z}\sim u^{4}$. For $A>5$, the near-boundary expansion of the field $\phi^{I}$ can then be organized in the following order:
$\phi^{I}(u,t,x^{i})\approx\left(0+\delta\phi_{(1)}^{I}(t,x^{i})\right)\,u^{5-A}+\dots+\left(\bar{\phi}_{(0)}^{I}(x^{i})+\delta\phi_{(0)}^{I}(t,x^{i})\right)+\mathcal{O}(u).$ (44)
Evidently, the background profile $\bar{\phi}^{I}=\bar{\phi}_{(0)}^{I}(x^{i})\sim x^{i}$ enters the subleading term in this expansion. It should be interpreted as a non-zero vacuum expectation value, $\langle\Phi^{I}\rangle\sim x^{i}$, related to a phase with SSB (in the standard quantization scheme). For the benchmark potential $V(X,Z)=X^{M}Z^{N}$, the condition $A>5$ requires that $M+2N>5/2$ Alberte:2017oqx . When $0<A<5$, the $u$-independent term dominates the UV expansion (43), and the profile $\bar{\phi}^{I}=\bar{\phi}_{(0)}^{I}(x^{i})\sim x^{i}$ is dual to the external source of the $\Phi^{I}$ operator that explicitly breaks translations.